A straight rod of length L = 20.2 cm, carries a uniform charge, ? = 2.85 x 10^-6 C/m. The rod is located along the y axis from y1= 0 to y2 = L. Find the expression or the electric field along the y axis, Ey, at point P. What is the magnitude of the field at yp = 55.0 cm?
Answer #1
Concepts and reason

Use the concept of electric field at a point due to a point charge.

Use the formula for the linear charge density to calculate the charge on the straight rod. Calculate the electric field at a point on the y-axis due to the small element of the rod, and finally integrate it to find the electric field due to entire straight rod at a point on the y-axis.

Fundamentals

The magnitude of electric field due to a point charge at a distance from the point charge is,

E=kqr2E = \frac{{kq}}{{{r^2}}}

Here, k is the Coulomb constant, q is the charge, and r is the distance.

The linear charge density is defined as the charge per unit length and is given by,

λ=qL\lambda = \frac{q}{L}

Here, q is the charge and L is the length.

In the following figure, a rod of length L is placed along the y-axis. Here, yp{y_p} is the distance of a point at which the electric field needs to be calculated.

y, = 55.0 cm

Consider an element of length dy at a distance y from the origin. The charge of the small element of length dydy is,

dq=λdydq = \lambda dy

From the figure, the distance of the element of charge dqdq from the point yp{y_p} is ypy{y_p} - y .

The electric field due to the small element dydy is,

dE=kdq(ypy)2dE = k\frac{{dq}}{{{{\left( {{y_p} - y} \right)}^2}}}

Here, ypy{y_p} - y is the distance between the small element of charge dqdq and the point yp{y_p} .

Substitute λdy\lambda dy for dqdq .

dE=kλdy(ypy)2dE = k\frac{{\lambda dy}}{{{{\left( {{y_p} - y} \right)}^2}}}

The total electric field due to the rod is,

dE=0Lkλdy(ypy)2E=kλ0Ldy(ypy)2=kλ(1ypy)0L\begin{array}{c}\\\int {dE} = \int\limits_0^L {k\frac{{\lambda dy}}{{{{\left( {{y_p} - y} \right)}^2}}}} \\\\E = k\lambda \int\limits_0^L {\frac{{dy}}{{{{\left( {{y_p} - y} \right)}^2}}}} \\\\ = k\lambda \left( {\frac{1}{{{y_p} - y}}} \right)_0^L\\\end{array}

Now,

E=kλ(1ypL1yp0)=kλ(ypyp+Lyp(ypL))=kλ(Lyp(ypL))\begin{array}{c}\\E = k\lambda \left( {\frac{1}{{{y_p} - L}} - \frac{1}{{{y_p} - 0}}} \right)\\\\ = k\lambda \left( {\frac{{{y_p} - {y_p} + L}}{{{y_p}\left( {{y_p} - L} \right)}}} \right)\\\\ = k\lambda \left( {\frac{L}{{{y_p}\left( {{y_p} - L} \right)}}} \right)\\\end{array}

Substitute 9.0×109Nm2/C29.0 \times {10^9}{\rm{ N}} \cdot {{\rm{m}}^2}/{{\rm{C}}^2} for k, 2.85×106C/m2.85 \times {10^{ - 6}}{\rm{ C/m}} for λ\lambda , 20.2 cm for L, and 55.0 cm for yp{y_p} .

E=(9.0×109Nm2/C2)(2.85×106C/m)20.2cm(1m100cm)(55.0cm(1m100cm))(55.0cm(1m100cm)20.2cm(1m100cm))=2.70×104N/C\begin{array}{c}\\E = \frac{{\left( {9.0 \times {{10}^9}{\rm{ N}} \cdot {{\rm{m}}^2}/{{\rm{C}}^2}} \right)\left( {2.85 \times {{10}^{ - 6}}{\rm{ C/m}}} \right)20.2{\rm{ cm}}\left( {\frac{{1{\rm{ m}}}}{{100{\rm{ cm}}}}} \right)}}{{\left( {55.0{\rm{ cm}}\left( {\frac{{1{\rm{ m}}}}{{100{\rm{ cm}}}}} \right)} \right)\left( {55.0{\rm{ cm}}\left( {\frac{{1{\rm{ m}}}}{{100{\rm{ cm}}}}} \right) - 20.2{\rm{ cm}}\left( {\frac{{1{\rm{ m}}}}{{100{\rm{ cm}}}}} \right)} \right)}}\\\\ = 2.70 \times {10^4}{\rm{ N/C}}\\\end{array}

Ans:

The magnitude of electric field due to the rod at a point yp=55.0cm{y_p} = 55.0{\rm{ cm}} is 2.70×104N/C2.70 \times {10^4}{\rm{ N/C}} .